How much greater is the area of a square that has one side equal to 4 feet than a rectangle that measures 3 feet long and 2 feet wide?
Answer by rfer(16322)

square
A=4^2
A=16
rectangle
A=3*2
A=6
difference
16-6=10 sq ft more in square